OpenStudy (anonymous):

Use Newton's method to approximate a root of the equation 5sin(x)=x as follows. Let x1=1 be the initial approximation. The second approximation x2 is -.8850033452 and the third approximation x3 is ? I got 2.377647049 but it was wrong

OpenStudy (anonymous):

f(x)=5sin(x) - x f'(x) = 5cos(x) - 1 x3 = -.8850033452 - f(-.885)/f'(-.885)

OpenStudy (anonymous):

-.8850033452 - (5*Sin[-.8850033452] + .8850033452)/(5* Cos[-.8850033452] - 1) = 0.492644

OpenStudy (anonymous):

thanks! ^-^

OpenStudy (anonymous):

i see what you did

OpenStudy (anonymous):

you did 1- f(x2)/f'(x2) instead of x2 - f(x2)/f'(x2)